# Sample-Efficient Constrained Reinforcement Learning with General Parameterization

### v1

@@ Abstract
We consider a constrained Markov Decision Problem (CMDP) where the goal of an agent is to maximize the expected discounted sum of rewards over an infinite horizon while ensuring that the expected discounted sum of costs exceeds a certain threshold. Building on the idea of momentum-based acceleration, we develop the Primal-Dual Accelerated Natural Policy Gradient (PD-ANPG) algorithm that guarantees an $\epsilon$ global optimality gap and $\epsilon$ constraint violation with $\mathcal{O}(\epsilon^{-3})$ sample complexity. This improves the state-of-the-art sample complexity in CMDP by a factor of $\mathcal{O}(\epsilon^{-1})$.
@@ 1 Introduction
In this article, we provide an affirmative answer to the above question. We propose a Primal-Dual-based Accelerated Natural Policy Gradient (PD-ANPG) algorithm to solve CMDPs with general parameterization. We theoretically prove that PD-ANPG achieves $\epsilon$ optimality gap and $\epsilon$ constraint violation with $\mathcal{O}(\epsilon^{-3})$ sample complexity (Theorem 1) that improves the SOTA $\tilde{\mathcal{O}}(\epsilon^{-4})$ sample complexity result of [4].
@@ 1.1 Challenges and Key Insights
Finally, Theorem 1 segregates the objective and constraint violation rates from Lagrange convergence. Corollary 1 shows that Lagrange convergence error is a decreasing function of $\zeta$ (the dual learning rate). One might, therefore, be tempted to make $\zeta$ arbitrarily small. However, our analysis shows that although small $\zeta$ leads to better objective convergence, it worsens the constraint violation rate. We demonstrate how to optimally choose $\zeta$ to reach the middle ground. It is the effect of $\zeta$ that deter us from reaching the optimal $\mathcal{O}(\epsilon^{-2})$ sample complexity.
@@ 1.2 Related Works
NPG-PD [9] $\mathcal{O}(\epsilon^{-2})$ Softmax
CRPO [10] $\mathcal{O}(\epsilon^{-4})$ Softmax
NPG-PD [9] $\mathcal{O}(\epsilon^{-6})$ General
CRPO [10] $\mathcal{O}(\epsilon^{-6})$ General
C-NPG-PDA [4] $\tilde{\mathcal{O}}(\epsilon^{-4})$ General
PD-ANPG (This Work) $\mathcal{O}(\epsilon^{-3})$ General
Lower Bound [11] $\Omega(\epsilon^{-2})$-
Table 1: Summary of sample complexity results on CMDP with parameterized policies.
@@
Constrained RL: The tabular setting is well investigated, and many model-based [28, 29, 30, 31] and model-free [10, 30, 32, 33] algorithms are available in the literature. In comparison, there are relatively fewer works on parameterized policies. Policy mirror descent-primal dual (PMD-PD) algorithm was proposed by [7] that achieves $\mathcal{O}(\epsilon^{-3})$ sample complexity for softmax policies. For the same parameterization, [8] achieved $\mathcal{O}(\epsilon^{-6})$ sample complexity via their proposed Online Primal-Dual Natural Actor-Critic Algorithm. The primal-dual Natural Policy Gradient algorithm suggested by [9] yields $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-6})$ sample complexities for softmax and general parametrization respectively. [10] also proposed a primal policy-based algorithm that works for both the softmax and general function approximation cases. The state of the art for the general function approximation is given by [4], where the $\tilde{\mathcal{O}}(\epsilon^{-4})$ sample complexity is obtained. This work improves upon this direction to obtain $\mathcal{O}(\epsilon^{-3})$ sample complexity. The comparisons are summarized in Table 1.
@@ 2 Formulation
Let us consider a constrained Markov Decision Process (CMDP) characterized by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,c,P,\gamma,\rho)$ where $\mathcal{S},\mathcal{A}$ denote the state space and the action space respectively, $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ defines the reward function, $c:\mathcal{S}\times\mathcal{A}\rightarrow[-1,1]$ is the cost function, $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ indicates the state transition kernel (where $\Delta(\mathcal{S})$ is the collection of all probability distributions over $\mathcal{S}$), $\gamma$ is the discount factor, and $\rho\in\Delta(\mathcal{S})$ is the initial state distribution. Note that both $\mathcal{S}$ and $\mathcal{A}$, in our setting, can potentially be compact sets of infinite size. However, for simplicity, we assume these sets to be countable. The range of the cost function is chosen to be $[-1,1]$, rather than $[0,1]$, to ensure that the constraint in our central optimization problem (defined later in (2)) is non-trivial. A policy, $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ is defined as a distribution over the action space for a given state of the environment. For a given policy, $\pi$, and a state-action pair $(s,a)$, we define the $Q$-value associated with $g\in\{r,c\}$ as follows.
@@ 3 Algorithm
$\begin{split}\omega^{*}_{\theta,\lambda}&\in{\arg\min}_{\omega\in\mathbb{R}^{\mathrm{d}}}L_{\nu^{\pi_{\theta}}_{\rho}}(\omega,\theta,\lambda),\\ \text{where}\leavevmode\nobreak\ L_{\nu^{\pi_{\theta}}_{\rho}}(\omega,\theta,\lambda)&\triangleq\dfrac{1}{2}\mathbf{E}_{(s,a)\sim\nu^{\pi_{\theta}}_{\rho}}\bigg{[}\dfrac{1}{1-\gamma}A^{\pi_{\theta}}_{\mathrm{L},\lambda}(s,a)-\omega^{\mathrm{T}}\nabla_{\theta}\log\pi_{\theta}(a|s)\bigg{]}^{2}\end{split}$ (9)
@@
Algorithm 2 has a nested loop structure. The outer loop runs $K$ number of times. At a given instance, $k$, of the outer loop, the policy parameter $\theta_{k}$ and the Lagrange parameter, $\lambda_{k}$ are updated via (20) and (21). The estimate, $\hat{J}_{c,\rho}(\theta_{k})$ is computed via Algorithm 19. On the other hand, $\omega_{k}$, the approximate value of the natural policy gradient, $\omega_{\theta_{k},\lambda_{k}}^{*}$, is obtained by iteratively minimizing $L_{\nu_{\rho}^{\pi_{\theta}}}(\cdot,\theta_{k},\lambda_{k})$ in $H$ number of inner loop steps via the Accelerated Stochastic Gradient Descent (ASGD) procedure as stated in [6]. ASGD comprises the iterative updates (15)--(18) with tunable learning parameters $(\alpha,\beta,\xi,\delta)$ followed by a tail-averaging step (19). The gradient estimate utilized in (16) and (18) is obtained via Algorithm 19. It is worth mentioning that existing NPG algorithms such as that given in [5] typically apply the SGD, rather than the ASGD procedure, to obtain $\omega_{k}$. The difference between these subroutines is that while SGD uses only the current gradient estimate to update $\omega_{k}$, ASGD considers the contribution of all previous gradient estimates (momentum) using its convoluted iteration and tail-averaging steps.
@@
6:         $\hat{G}\leftarrow\hat{\nabla}_{\omega}L_{\nu^{\pi_{\theta}}_{\rho}}(\omega,\theta_{k},\lambda_{k})\big{|}_{\omega=\mathbf{y}_{h}}$ (Algorithm 19)
@@ 4 Analysis
The score function (written below) is $G$-Lipschitz and $B$-smooth where $B,G>0$. Formally, the following relations hold $\forall\theta,\theta_{1},\theta_{2}\in\Theta$, and $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$.
@@
Assumption 2 implies the following properties of the Lagrange function.
@@
If Assumption 2 holds, then the following inequalities are true $\forall\theta\in\Theta$ and $\forall\lambda\in\Lambda$.
$(a)\leavevmode\nobreak\ \|\nabla_{\theta}J_{\mathrm{L},\rho}(\theta,\lambda)\|\leq\dfrac{G(1+\lambda)}{(1-\gamma)^{2}},\leavevmode\nobreak\ \text{and}$
$(b)\leavevmode\nobreak\ J_{\mathrm{L},\rho}(\cdot,\lambda)\leavevmode\nobreak\ \text{is}\leavevmode\nobreak\ L(1+\lambda)-\text{smooth where}\leavevmode\nobreak\ L\triangleq\dfrac{B}{(1-\gamma)^{2}}+\dfrac{2G^{2}}{(1-\gamma)^{3}}$
@@
Statement (a) can be proven using (8) along with Assumption 2 and the fact that $|A_{\mathrm{L},\lambda}^{\pi_{\theta}}(s,a)|$ is bounded by $(1+\lambda)/(1-\gamma)^{2}$. Statement (b) follows from Proposition 4.2 of [36]. ∎
Lemma 2 establishes that Assumption 2 enforces the Lagrange function to be smooth and Lipschitz continuous with respect to its first argument. This property will turn out to be pivotal in our analysis.
@@ 4.1 Local-to-Global Convergence Lemma
$\begin{split}\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\|\omega_{k}\|^{2}&=\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\|\omega_{k}-\omega_{k}^{*}\|^{2}+\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\|\omega_{k}^{*}\|^{2}\\ &\overset{(a)}{\leq}\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\|\omega_{k}-\omega_{k}^{*}\|^{2}+\dfrac{1}{\mu_{F}^{2}K}\sum_{k=0}^{K-1}\mathbf{E}\left\|\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\right\|^{2}\end{split}$ (23)
where (a) utilises $\omega_{k}^{*}=F_{\rho}(\theta_{k})^{\dagger}\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})$ and Assumption 4. The second term in (23) can be further bounded by Lemma 4 (described below).
Lemma 4.
The following relation holds for $\eta=\frac{\mu_{F}^{2}}{4G^{2}L(1+\lambda_{\max})}$ where $L$ is given in Lemma 2 and $\{\theta_{k},\lambda_{k}\}_{k=0}^{K-1}$ are recursively obtained via (20) and (21).
$\begin{split}\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\|\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\|^{2}&\leq(2G^{4}+\mu_{F}^{2})\left(\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\|\omega_{k}-\omega_{k}^{*}\|^{2}\right)\\ &+\left(\dfrac{32G^{4}L(1+\lambda_{\max})^{2}}{\mu_{F}^{2}(1-\gamma)}\right)\dfrac{1}{K}+\left(\dfrac{16G^{4}L(1+\lambda_{\max})}{\mu_{F}^{2}(1-\gamma)^{2}}\right)\zeta\end{split}$ (24)
As a result of (23) and (24), the term $\frac{1}{K}\sum_{0\leq k<K}\mathbf{E}\|\omega_{k}\|^{2}$ can now be bounded by a term proportional to $\frac{1}{K}\sum_{0\leq k<K}\mathbf{E}\|\omega_{k}-\omega_{k}^{*}\|^{2}$ plus some additional functions of $K$ and the Lagrange update parameter, $\zeta$. This indicates that to ensure global convergence, $\zeta$ must also be judiciously chosen as a decreasing function of the horizon, $K$. Our next subsection provides a bound on $\mathbf{E}\|\omega_{k}-\omega_{k}^{*}\|^{2}$ and the first order term $\mathbf{E}\|\mathbf{E}[\omega_{k}|\theta_{k},\lambda_{k}]-\omega_{k}^{*}\|$.
@@ 4.2 Local Convergence of the Natural Policy Gradient
$\begin{split}&\bar{\mathbf{y}}_{h}=\alpha\bar{\mathbf{x}}_{h}+(1-\alpha)\bar{\mathbf{v}}_{h},\leavevmode\nobreak\ \bar{\mathbf{x}}_{h+1}=\bar{\mathbf{y}}_{h}-\delta\nabla_{\omega}L_{\nu^{\pi_{\theta}}_{\rho}}(\omega,\theta,\lambda)\\ &\bar{\mathbf{z}}_{h}=\beta\bar{\mathbf{y}}_{h}+(1-\beta)\bar{\mathbf{v}}_{h},\leavevmode\nobreak\ \bar{\mathbf{v}}_{h+1}=\bar{\mathbf{z}}_{h}-\xi\nabla_{\omega}L_{\nu^{\pi_{\theta}}_{\rho}}(\omega,\theta,\lambda)\end{split}$ (28)

### v3

@@ Abstract
We consider a constrained Markov Decision Problem (CMDP) where the goal of an agent is to maximize the expected discounted sum of rewards over an infinite horizon while ensuring that the expected discounted sum of costs exceeds a certain threshold. Building on the idea of momentum-based acceleration, we develop the Primal-Dual Accelerated Natural Policy Gradient (PD-ANPG) algorithm that ensures an $\epsilon$ global optimality gap and $\epsilon$ constraint violation with $\tilde{\mathcal{O}}((1-\gamma)^{-7}\epsilon^{-2})$ sample complexity for general parameterized policies where $\gamma$ denotes the discount factor. This improves the state-of-the-art sample complexity in general parameterized CMDPs by a factor of $\mathcal{O}((1-\gamma)^{-1}\epsilon^{-2})$ and achieves the theoretical lower bound in $\epsilon^{-1}$.
@@ 1 Introduction
In this article, we provide an affirmative answer to the above question. We propose a Primal-Dual-based Accelerated Natural Policy Gradient (PD-ANPG) algorithm to solve $\gamma$-discounted CMDPs with general parameterization. We theoretically prove that PD-ANPG achieves $\epsilon$ optimality gap and $\epsilon$ constraint violation with $\tilde{\mathcal{O}}((1-\gamma)^{-7}\epsilon^{-2})$ sample complexity (Theorem 1) that improves the SOTA $\tilde{\mathcal{O}}((1-\gamma)^{-8}\epsilon^{-4})$ sample complexity result of [4]. It closes the gap between the theoretical upper and lower bounds of sample complexity in general parameterized CMDPs (in terms of $\epsilon^{-1}$), which was an open problem for quite some time (see the results of [5], [6], [4] in Table 1).
@@ 1.1 Challenges and Key Insights
Finally, Theorem 1 segregates the objective and constraint violation rates from Lagrange convergence. Corollary 1 shows that Lagrange convergence error is bounded by an independent function of $\zeta$ (the dual learning rate). One might, hence, be tempted to make $\zeta$ arbitrarily small. However, our analysis shows that although small $\zeta$ leads to better objective convergence, it worsens the constraint violation rate. We demonstrate how to optimally choose $\zeta$ to reach the middle ground which eventually leads us to $\tilde{\mathcal{O}}(\epsilon^{-2})$ sample complexity.
@@ 1.2 Related Works
NPG-PD [5] $\mathcal{O}((1-\gamma)^{-5}\epsilon^{-2})$ Softmax
CRPO [6] $\mathcal{O}((1-\gamma)^{-7}\epsilon^{-4})$ Softmax
NPG-PD [5] $\mathcal{O}((1-\gamma)^{-8}\epsilon^{-6})$ General
CRPO [6] $\mathcal{O}((1-\gamma)^{-13}\epsilon^{-6})$ General
C-NPG-PDA [4] $\tilde{\mathcal{O}}((1-\gamma)^{-8}\epsilon^{-4})$ [^1] We would like to point out that the sample complexity of C-NPG-PDA reported in [4] is $\tilde{\mathcal{O}}((1-\gamma)^{-6}\epsilon^{-4})$. However, the result is erroneous. The authors have subsequently corrected their result, and the sample complexity has been modified to $\tilde{\mathcal{O}}((1-\gamma)^{-8}\epsilon^{-4})$ in the arXiv version (updated May 2024). General
PD-ANPG (This Work) $\tilde{\mathcal{O}}((1-\gamma)^{-7}\epsilon^{-2})$ General
Lower Bound [11] $\Omega((1-\gamma)^{-5}\epsilon^{-2})$-
Table 1: Summary of sample complexity results on CMDP with parameterized policies. The parameter $\gamma$ indicates the discount factor. The dependence of the sample complexities of PMD-PD and PD-NAC on $\gamma$ is not depicted in [9, 10].
@@
Constrained RL: The tabular setting is well investigated, and many model-based [28, 29, 30, 31] and model-free [6, 30, 32, 33] algorithms are available in the literature. In comparison, there are relatively fewer works on parameterized policies. Policy mirror descent-primal dual (PMD-PD) algorithm was proposed by [9] that achieves $\mathcal{O}(\epsilon^{-3})$ sample complexity for softmax policies. For the same parameterization, [10] achieved $\mathcal{O}(\epsilon^{-6})$ sample complexity via their proposed Online Primal-Dual Natural Actor-Critic Algorithm. The primal-dual Natural Policy Gradient algorithm suggested by [5] yields $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-6})$ sample complexities for softmax and general parametrization respectively. [6] also proposed a primal policy-based algorithm that works for both the softmax and general function approximation cases. The state of the art for the general parameterization is given by [4], where the $\tilde{\mathcal{O}}(\epsilon^{-4})$ sample complexity is obtained. This work improves upon this direction to obtain $\tilde{\mathcal{O}}(\epsilon^{-2})$ sample complexity. The comparisons are summarized in Table 1.
@@ 2 Formulation
Let us consider a constrained Markov Decision Process (CMDP) characterized by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,c,P,\gamma,\rho)$ where $\mathcal{S},\mathcal{A}$ denote the state space and the action space respectively, $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ defines the reward function, $c:\mathcal{S}\times\mathcal{A}\rightarrow[-1,1]$ is the cost function, $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ indicates the state transition kernel (where $\Delta(\mathcal{S})$ is the collection of all probability distributions over $\mathcal{S}$), $\gamma$ is the discount factor, and $\rho\in\Delta(\mathcal{S})$ is the initial state distribution. Note that the state space $\mathcal{S}$, in our setting, can potentially be a compact set of infinite size. However, for simplicity, we assume it to be countable. The action space, $\mathcal{A}$, is assumed to be of finite size. The range of the cost function is chosen to be $[-1,1]$, rather than $[0,1]$, to ensure that the constraint in our central optimization problem (defined later in (2)) is non-trivial. A policy, $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ is defined as a distribution over the action space for a given state of the environment. For a given policy, $\pi$, and a state-action pair $(s,a)$, we define the $Q$-value associated with $g\in\{r,c\}$ as follows.
@@ 3 Algorithm
$\begin{split}\omega^{*}_{\theta,\lambda}&\in{\arg\min}_{\omega\in\mathbb{R}^{\mathrm{d}}}L_{\nu^{\pi_{\theta}}_{\rho}}(\omega,\theta,\lambda),\\ \text{where}\leavevmode\nobreak\ L_{\nu^{\pi_{\theta}}_{\rho}}(\omega,\theta,\lambda)&\triangleq\dfrac{1}{2}\mathbf{E}_{(s,a)\sim\nu^{\pi_{\theta}}_{\rho}}\left[\left(\dfrac{1}{1-\gamma}A^{\pi_{\theta}}_{\mathrm{L},\lambda}(s,a)-\omega^{\mathrm{T}}\nabla_{\theta}\log\pi_{\theta}(a|s)\right)^{2}\right]\end{split}$ (9)
@@
Algorithm 2 has a nested loop structure. The outer loop runs $K$ number of times. At a given instance, $k$, of the outer loop, the policy parameter $\theta_{k}$ and the Lagrange parameter, $\lambda_{k}$ are updated via (20) and (21). The estimate, $\hat{J}_{c,\rho}(\theta_{k})$ is computed via Algorithm 19. On the other hand, $\omega_{k}$, the approximate value of the natural policy gradient $\omega_{\theta_{k},\lambda_{k}}^{*}$ is obtained by iteratively minimizing $L_{\nu_{\rho}^{\pi_{\theta_{k}}}}(\cdot,\theta_{k},\lambda_{k})$ in $H$ number of inner loop steps via the Accelerated Stochastic Gradient Descent (ASGD) procedure as stated in [8]. ASGD comprises the iterative updates (15)--(18) with tunable learning parameters $(\alpha,\beta,\xi,\delta)$ followed by a tail-averaging step (19). The gradient estimate utilized in (16) and (18) is obtained via Algorithm 19. It is worth mentioning that existing NPG algorithms such as that given in [7] typically apply the SGD, rather than the ASGD procedure, to obtain $\omega_{k}$. The difference between these subroutines is that while SGD uses only the current gradient estimate to update $\omega_{k}$, ASGD considers the contribution of all previous gradient estimates (momentum) using its convoluted iteration and tail-averaging steps.
@@
6:         $\hat{G}\leftarrow\hat{\nabla}_{\omega}L_{\nu^{\pi_{\theta_{k}}}_{\rho}}(\omega,\theta_{k},\lambda_{k})\big{|}_{\omega=\mathbf{y}_{h}}$ (Algorithm 19)
@@ 4 Analysis
The log-likelihood function is $G$-Lipschitz and $B$-smooth where $B,G>0$. Formally, the following relations hold $\forall\theta,\theta_{1},\theta_{2}\in\Theta$, and $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$.
@@
Assumption 2 implies the boundedness of the gradient of the Lagrange function. This can be formally expressed as follows.
@@
If Assumption 2 holds, then the following inequality is true $\forall\theta\in\Theta$ and $\forall\lambda\in\Lambda$.
$\|\nabla_{\theta}J_{\mathrm{L},\rho}(\theta,\lambda)\|\leq\dfrac{G(1+\lambda_{\max})}{(1-\gamma)^{2}}$
@@
Statement (a) can be proven using (8) along with Assumption 2 and the facts that $|A_{\mathrm{L},\lambda}^{\pi_{\theta}}(s,a)|$ is bounded by $(1+\lambda)/(1-\gamma)$ and $\lambda\leq\lambda_{\max}$, $\forall\lambda\in\Lambda$. ∎
The result established by Lemma 2 will be pivotal in our further analysis.
@@ 4.1 Local-to-Global Convergence Lemma
$\begin{split}\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\|\omega_{k}\|^{2}&\leq\dfrac{2}{K}\sum_{k=0}^{K-1}\mathbf{E}\|\omega_{k}-\omega_{k}^{*}\|^{2}+\dfrac{2}{K}\sum_{k=0}^{K-1}\mathbf{E}\|\omega_{k}^{*}\|^{2}\\ &\overset{(a)}{\leq}\dfrac{2}{K}\sum_{k=0}^{K-1}\mathbf{E}\|\omega_{k}-\omega_{k}^{*}\|^{2}+\dfrac{2}{\mu_{F}^{2}K}\sum_{k=0}^{K-1}\mathbf{E}\left\|\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\right\|^{2}\\ &\overset{(b)}{\leq}\dfrac{2}{K}\sum_{k=0}^{K-1}\mathbf{E}\|\omega_{k}-\omega_{k}^{*}\|^{2}+\dfrac{2G^{2}(1+\lambda_{\max})^{2}}{\mu_{F}^{2}(1-\gamma)^{4}}\end{split}$ (23)
where (a) utilises $\omega_{k}^{*}=F_{\rho}(\theta_{k})^{\dagger}\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})$ and Assumption 4. The second inequality applies Lemma 2 together with the fact that $\lambda_{k}\in\Lambda$. Our next subsection provides a bound on $\mathbf{E}\|\omega_{k}-\omega_{k}^{*}\|^{2}$ and the first order term $\mathbf{E}\|\mathbf{E}[\omega_{k}|\theta_{k},\lambda_{k}]-\omega_{k}^{*}\|$.
@@ 4.2 Local Convergence of the Natural Policy Gradient
$\begin{split}&\bar{\mathbf{y}}_{h}=\alpha\bar{\mathbf{x}}_{h}+(1-\alpha)\bar{\mathbf{v}}_{h},\\ &\bar{\mathbf{x}}_{h+1}=\bar{\mathbf{y}}_{h}-\delta\nabla_{\omega}L_{\nu^{\pi_{\theta_{k}}}_{\rho}}(\omega,\theta_{k},\lambda_{k})\big{|}_{\omega=\bar{\mathbf{y}}_{h}}\\ &\bar{\mathbf{z}}_{h}=\beta\bar{\mathbf{y}}_{h}+(1-\beta)\bar{\mathbf{v}}_{h},\\ &\bar{\mathbf{v}}_{h+1}=\bar{\mathbf{z}}_{h}-\xi\nabla_{\omega}L_{\nu^{\pi_{\theta_{k}}}_{\rho}}(\omega,\theta_{k},\lambda_{k})\big{|}_{\omega=\bar{\mathbf{y}}_{h}}\end{split}$ (27)